# SOOD++: Leveraging Unlabeled Data to Boost Oriented Object Detection

### v1

@@ 7 Conclusion
This paper proposes a practical solution named SOOD++ for semi-supervised oriented object detection. Focusing on oriented objects’ characteristics in aerial scenes, we customize Geometry-aware Adaptive Weighting (GAW) loss and Noise-driven Global Consistency (NGC). The former considers the importance of geometry information for oriented objects. The latter introduces the global layout concept to SSOD, measuring the global similarity between the teacher and the student in a many-to-many manner. Additionally, a Simple Instance-aware Dense Sampling (SIDS) strategy is proposed to construct high-quality pseudo-labels. To validate the effectiveness of our method, we have conducted extensive experiments on several challenging datasets. SOOD++ achieves consistent improvements under partially and fully labeled data settings, compared with SOTA methods. We hope this work will encourage the community to pay attention to semi-supervised oriented object detection and facilitate future research.

### v2

@@ 6 Conclusion
In this paper, we present SOOD++, a simple yet solid framework for semi-supervised oriented object detection (SSOD) tailored to aerial scenes. The integration of the sampling strategy, intrinsic geometric priors, and global consistency in SOOD++ offers a novel approach that significantly enhances detection performance by mining useful information from massive unlabeled data. Additionally, this approach provides insights that could inspire further research in both semi-supervised and oriented object detection fields. To validate the effectiveness of our method, we have conducted extensive experiments on six challenging datasets. Compared with SOTA methods, SOOD++ achieves consistent improvements under partially and fully labeled data settings. We hope SOOD++ could be a foundation for future advancements in oriented object detection, particularly in semi-supervised learning.
Limitation and future work. Despite the significant progress made, several important properties of oriented objects, such as occlusions, and texture, remain largely unexplored in our work. Incorporating these factors could further enhance the model’s performance. Additionally, oriented objects and complex objects also appear in other tasks, including 3D object detection and text detection, presenting an interesting direction for future exploration.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants 62225603, U2341227, and 623B2038. The authors thank Wei Xu, Xin Zhou, Xingkui Zhu, and Xingyu Jiang for their constructive comments on earlier drafts of this manuscript.